A 32Okg wooden raft floats on a lake. When a 75kg man stands on the raft, it sinks 3.5cm deeper into the water. When he steps off, the raft oscillates for a while.


i. What is the frequency of oscillation? 


ii. What is the total energy of oscillation (ignoring damping)? 


1
Expert's answer
2021-09-29T17:49:59-0400

The additional buoyancy force when the man is standing:


"F_a=\rho g\Delta V=\rho g(A\Delta x)."

The angular frequency of oscillations can be expressed as


"a=\omega^2\Delta x,\\\space\\
\omega=\sqrt\frac a{\Delta x}."

The restoring force that appears when the man is standing is, according to Newton's second law:


"F_r=F_a,\\
F_r=Ma,\\\space\\
a=\frac{F_r}{M}=\frac{F_a}{M},\\\space\\
\omega=\sqrt{\frac{F_a}{M\Delta x}}=\sqrt{\frac{\rho gA\Delta x}{M\Delta x}}=\sqrt{\frac{\rho gA}{M}}."

Find the area from the equilibrium of forces for the wood floating alone and with the man standing on it:


"Mg=\rho gAh,\\
(M+m)g=\rho gA(h+\Delta x),"

divide one by another:


"\frac{Mg}{(M+m)g-\rho gA\Delta x}=1,\\\space\\
\rho gA\Delta x=mg,\\\space\\
\rho gA=\frac{mg}{\Delta x}."

Finally,


"f=\frac\omega{2\pi}=\frac{1}{2\pi}\sqrt{\frac{\rho gA}{M}}=\frac{1}{2\pi}\sqrt{\frac{mg}{M\Delta x}}\\\space\\
f=\frac{1}{2\pi}\sqrt{\frac{75·9.8}{320·0.035}}=1.29\text{ Hz}."

The total energy is


"E=\frac 12Mv^2=\frac12 M(\omega\Delta x)^2=\frac12 M(2\pi f\Delta x)^2=\\=12.9\text{ J}."